OpenStudy (anonymous):

-4-7/4x=10 Solve the following equation. Please show all work! Have done just want a second opinion

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

x=14

OpenStudy (phi):

is the problem \[ -4 - \frac{7}{4x} = 10\]

OpenStudy (phi):

if so, \[-4 - \frac{7}{4\cdot 14}=10\] which does not work. if the problem is \[ -4 - \frac{7}{4}x=10 \] then with x=14 \[ -4 - \frac{7}{4}14=10 \] which also does not work.

OpenStudy (anonymous):

I added 4 to both sides

OpenStudy (phi):

that is a good start. but which problem is it (see above)?

OpenStudy (anonymous):

I subtracted the ricp from it too, I guess I just need to start over again

OpenStudy (phi):

Let's do -4 - (7/4)x = 10 (the x is up top) then add 4 to both sides -(7/4)x= 14 this is what you got, right? to get rid of the -7/4 *multiply* by the reciprocal on both sides of the equation

OpenStudy (anonymous):

yes

OpenStudy (phi):

can you do \[ \frac{-4}{7}\cdot \frac{-7}{4}x= 10\cdot \frac{-4}{7} \]

OpenStudy (anonymous):

cant cancel can you since the 4 and 7 are oposit of each other

OpenStudy (anonymous):

-4/7*-7/4 =0

OpenStudy (phi):

You can cancel (that is why we are doing it). the order of multiplication does not matter 2*3= 3*2 so \[ \frac{-4}{7}\cdot\frac{-7}{4} =\frac{-7}{7}\cdot\frac{-4}{4} = -1\cdot -1= 1\]

OpenStudy (anonymous):

x=-4/70?

OpenStudy (anonymous):

just makes it x

OpenStudy (phi):

close, but think of the 10* -4/7 as \[ \frac{10}{1}\cdot \frac{-4}{7} \] multiply top times top and bottom times bottom

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

-4/70

OpenStudy (phi):

Let's back up: -4 - (7/4)x = 10 then add 4 to both sides -(7/4)x = 14 x = 14 * -4/7 can you finish this?

OpenStudy (anonymous):

56/7

OpenStudy (phi):

what happen to the minus sign?

OpenStudy (anonymous):

forgot to put it in

OpenStudy (anonymous):

-56/7

OpenStudy (phi):

OK, let's check your answer x= -56/7 you start with -4 - (7/4)x = 10 if x is -56/7 then the left-hand side is -4 - (7/4)*(-56/7) can you do this?

OpenStudy (phi):

before you forget notice that it is minus times minus so -7/4 * -56/7 is + (cancel the minus signs) also, you can cancel the 7's

OpenStudy (anonymous):

14

OpenStudy (phi):

keep going, you are doing -4 - (7/4)*(-56/7)

OpenStudy (phi):

you found -7/4 * -56/7 = +14 so now -4 + 14 =

OpenStudy (anonymous):

-4(14)=56

OpenStudy (anonymous):

-56

OpenStudy (phi):

look at it more carefully we are adding

OpenStudy (anonymous):

18

OpenStudy (phi):

-4 - (7/4)*(-56/7) --> -4 + 14

OpenStudy (phi):

4+14= 18 -4 + 14 = ?

OpenStudy (phi):

when you add negative numbers, it's like subtraction.

OpenStudy (anonymous):

-10?

OpenStudy (phi):

-4 + 14 is the same problem as 14 - 4

OpenStudy (phi):

Here is a 5 minute video that might help http://www.khanacademy.org/video/adding-negative-numbers?playlist=Developmental+Math+2

OpenStudy (anonymous):

Thank you it might I hope so thanks for all of your help

OpenStudy (phi):

you will get +10 which matches the original -4 - (7/4)x = 10 you figured out that when x= -56/7 that the equation turns into 10 = 10 which means you got the right answer.

OpenStudy (anonymous):

With lots of help from you thanks

OpenStudy (phi):

If the video helps and is not too boring, then search for the videos on solving equations (use his search window). If something does not make sense, search for a video that explains it, because he explains *everything*

OpenStudy (anonymous):

Thank you I will look at the links wheather they are boring or not I need the info